Im Very Confused on how to find x and y intercepts , i try my best to work them out but i need help with graphing them. I cannot understand how to graph 7x-3y=6
Answer by Alan3354(69443)

Im Very Confused on how to find x and y intercepts , i try my best to work them out but i need help with graphing them. I cannot understand how to graph 7x-3y=6
-------------------
At the y-intercept, x = 0
7x- 3y = 6
7*0 - 3y = 6
y = -2
--> (0,-2)
=======================
At the x-intercept, y = 0
7x- 3y = 6
7x - 3*0 = 6
x = 6/7
--> (6/7,0)
--------------
You can plot those 2 points and draw a straight line thru them.